# A microscopic theory of Anderson localization of electrons in random lattices

### v1

@@ Abstract
The existence of Anderson localization, characterized by vanishing diffusion due to strong randomness, has been demonstrated in numerous ways. A systematic approach based on the Anderson quantum model of the Fermi gas in random lattices that can describe both diffusive and localized regimes has not yet been fully established. We build upon a recent publication [1] and present a microscopic theory of disordered electrons covering both the metallic phase with extended Bloch waves and the localized phase where the propagating particle forms a quantum bound state with the hole left behind at the origin. The general theory provides a framework for constructing controlled approximations to one-particle and two-particle Green functions that satisfy the necessary conservation laws and causality requirements in the whole range of disorder strength. It is used explicitly to derive a local, mean-field-like approximation for the two-particle irreducible vertices, enabling quantitative analysis of the solution’s properties in both metallic and localized phases, including critical behavior at the Anderson localization transition.
@@ I Introduction
Anderson formulated a simple model of the propagation of a quantum particle in a random lattice [2]. His solution, in which diffusion is suppressed over long distances when disorder is sufficiently strong, known as Anderson localization, sparked tremendous interest in the physics community. Early research concentrated on finding the mobility edge between the extended and localized states of the Fermi gas in models of disordered alloys [3, 4, 5, 6, 7, 8]. The vanishing of diffusion was attributed to zero of the static electric conductivity. Based on qualitative considerations and experimental evidence, Mott suggested a scenario of a discontinuous metal-insulator transition at the mobility edge with a minimal metallic conductivity [5, 9]. Symmetry arguments and the apparent short-range behavior of the averaged resolvent led to a conjecture that the critical behavior of interacting matrices, nonlinear $\sigma$-model, may simulate the mobility edge behavior in dimensions $2+\epsilon$ [10, 11, 12].
@@
The analogy between the Anderson localization transition and the critical behavior of phase transitions in statistical mechanics was supported by scaling arguments in Ref. [13]. Since then, the focus has shifted from the original Anderson quantum model of disordered non-interacting electrons to classical statistical mechanical models, allowing the use of established techniques for studying the critical behavior of interacting systems, such as scaling properties and universality classes [14]. The models differed in the universality class [11] or the underlying lattice [7]. In particular, a solution on the Bethe lattice with infinite coordination number was considered as a mean-field solution of the Anderson localization transition. Unlike the regular lattices, the critical behavior at the localization transition was found to be exponential [15, 16]. The exponential critical behavior was later shown to be an artifact of the Bethe lattice [17]. Moreover, the man-field solution on the Bethe lattice was later shown to miss the correct high-dimensional limit of the Cayley tree [18, 19]. The Bethe lattice is the infinite limit of a regular random graph [20], and the concept of Anderson localization was then extended to random graphs [21, 22, 23, 24].
@@
The behavior of random matrix models in low spatial dimensions, $d\leq 2$, was known to have a localized spectrum [25]. The existence of the Anderson localization transition and the mobility edge in dimensions $d>2$ was proved rigorously [26]. Numerical simulations in $d=3$ complemented the mathematical proof [27, 28]. The concept of Anderson localization has grown beyond its original formulation of charge diffusion and has been extended to light propagation [29, 30], the localization of thermal phonons [31], and the Anderson transition in cold atoms [32].
@@
All these approaches focused on the critical behavior of the Anderson localization in the insulating phase. It means they use either localized states or a diverging scale, being the localization length. They are inapplicable in the metallic phase with no evident diverging scale. Instead, the limit of the static conductivity to zero has been used to indicate the Anderson localization transition. To find a theory of the Anderson localization transition approached from the metallic phase, the diagrammatic expansion in the Anderson model of disordered electrons, together with Green functions, was used to derive a self-consistent theory for the diffusion constant that turns zero at the localization transition [33, 34, 35, 36, 37]. It has been assumed to be an effective mean-field-like theory, for which, however, systematic extensions have not yet been found.
@@
Over the last six decades, we have gained significant insights into the Anderson localization transition and its associated critical behavior in the localized phase, characterized by a diverging localization length. However, not all questions about the origin of the localized phase have been answered. First, the critical scale is defined only in the localized phase. Second, no order parameter analogous to those in statistical mechanics has been identified. Third, the microscopic origin of the localized states, as described by the Anderson model, remains unexplained in the existing approaches.
@@
The way to a microscopic theory of Anderson localization is to use the original model of a Fermi gas in a random lattice and attempt to solve it in a controlled manner. It is provided by the diagrammatic expansion of the random potential in the averaged one-particle and two-particle resolvents. The well-established local mean-field theory for the electronic properties of random alloys, the coherent-potential approximation (CPA) [38, 39], does not lead to Anderson localization. A controlled way to go beyond the local approximation is to use the limit where the spatial dimension tends to infinity. The first fully self-consistent perturbation theory for single-electron properties, exact in infinite spatial dimensions, was shown to be equivalent to CPA [40, 41]. It misses vertex corrections to the Drude conductivity and is unfit to describe Anderson localization [42, 43]. Vertex corrections diminishing the electric conductivity are caused by backscattering and non-local correlations missed by the CPA [44, 45].
@@
The charge diffusion and its vanishing depend on how far a particle can propagate in a random environment. This information is not contained in the one-particle self-energy. It is the electron-hole correlation function, derived from the two-particle Green’s function, that carries information about the origin of the electron propagation. It is determined from a Bethe-Salpeter equation with a two-particle irreducible vertex. The Anderson localization transition is signalled by a bifurcation point at which a new solution for the two-particle vertex emerges. Self-consistent nonlinear equations for the two-particle vertex must be used to see the bifurcation in the two-particle vertex [46]. We earlier used the parquet construction combining Bethe-Salpeter equations for nonlocal irreducible vertices in the electron-hole and electron-electron scattering channels to introduce a two-particle self-consistency [47, 48]. This approach is similar to Refs. [33, 34, 37], who replaced the full two-particle self-consistency of the parquet equations for the irreducible vertices by a self-consistent equation for the dynamical diffusion coefficient. We identified a critical point that resembles the Anderson localization transition. Our solution, however, seemed unphysical since it did not obey the necessary Ward identity between the self-energy and the irreducible electron-hole vertex of Ref. [34]. This Ward identity guarantees the correct low-energy asymptotics of the diffusion pole in the density response function of the metallic phase. The dynamical vertex in the perturbation expansion is causal so that the related self-energy has the demanded analytic properties. We showed that the microscopic causality is incompatible with the macroscopic Ward identity [49, 50]. The observed macroscopic quantities, the electron-hole correlation function and conductivity, must conform to the conservation laws. We showed how a conserving vertex complying with the desired Ward identity is obtained from the dynamical one of the perturbation theory in Ref. [51]. This addition enabled us to establish a controlled method for relating the outcome of microscopic perturbation theory to macroscopic observables. It opened a route towards a consistent microscopic theory of Anderson localization.
@@
I made the first step towards a fully controllable microscopic theory of Anderson localization in a recent publication [1]. A local approximation of the two-particle irreducible vertex in the parquet equations led to a critical point in the metallic phase of the Anderson model of disordered electrons. The critical point was identified with the Anderson localization transition. The diverging scale at the transition point was identified as the frequency derivative of the dynamical electron-hole irreducible vertex, or the dynamical conductivity. It was found to match the divergence of the localization length at the other side of the Anderson localization transition. The most fundamental conclusion of this paper was that the localized state is a quantum bound state of the propagating electron and the hole left at the beginning of its motion.
@@
This paper aims to build upon Ref. [1] and to present a fully controllable microscopic theory of Anderson localization based on self-consistent non-linear equations for two-particle irreducible vertices resulting from the parquet theory interconnecting multiple scattering in the electron-hole, diffuson, and electron-electron, cooperon, channels. We show that, to reach the critical point signaling a transition to the localized phase, two-particle functions must contain more information than that provided by the self-energy. The localized phase contains quantum bound states characterized by a new order parameter that induces a gap in the two-particle vertex function. Since there is no condensation of the bound states, no gap in the one-particle self-energy emerges. We present details of the localized bound states, determine the threshold energy for bond breaking, and turn the localized state metallic. We also discuss the relationship between the microscopic dynamical quantities derived from perturbation theory and the macroscopic observables that obey conservation laws. We further specify the extent to which the Ward identities are valid in the metallic and localized phases. We demonstrate that linear-response theory breaks down in the localized phase and that static theories cannot describe Anderson localization. At the end, we discuss the relationship between the dynamical conductivity, the electron-hole correlation function, and the complex diffusion function in both metallic and localized phases.
@@ II.1 Hamiltonian and averaged resolvent
We assume the validity of the ergodic hypothesis, which states that spatial averaging is equivalent to configurational averaging. Consequently, all physical quantities in the thermodynamic limit will be configurationally averaged and translationally invariant. The averaged functions of interest, which determine all relevant physical quantities, are the one-particle and two-particle resolvents in the complex energy plane. The matrix representation of the one-particle resolvent for complex energy $z=E+i\eta$ in the direct space reads
@@
where $\epsilon(\mathbf{k})$ is the lattice dispersion relation. The second equality is the Dyson equation, which introduces the self-energy, $\Sigma_{\mathbf{k}}(z)$, containing the entire information about the impact of the random potential on the electron’s propagation. The averaged one-particle Green function determines the electronic structure and the spectral properties of the Fermi gas in a random lattice. It contains no information about the extent of the configurationally dependent electron wave functions at the given energy, $z=E+i0^{+}$.
@@ II.2 Two-particle vertex and electron-hole correlation function
The extent of the eigenstates of the random Hamiltonian affects the two-particle resolvent or the averaged two-particle Green function. The two-particle Green function of non-interacting particles in a random potential has generally two complex energies $z_{1}$ and $z_{2}$
@@
The behavior of the two-particle Green function with energies $z_{1}\equiv E_{+}=E+\omega/2+i0^{+}$ and $z_{2}\equiv E_{-}=E-\omega/2-i0^{+}$ is decisive for the determination whether the eigenstates of the random Hamiltonian are localized or extended. Here $E$ is the Fermi energy and $\omega$ is the transfer energy between particle and hole. The two energies characterize the electron-hole resolvent $\mathcal{G}^{RA}$, the wave-vector representation of which is
@@
with the Fermi energy $E$. The transfer energy $\omega$ and wave vector $\mathbf{q}$ are conserved during the scattering processes when decomposing the two-particle Green function into elementary scattering events. The dynamical variables affected by the random potential are wave vectors $\mathbf{k},\mathbf{k}^{\prime}$.
The averaged two-particle Green function $\mathcal{G}^{RA}_{{\bf k}{\bf k}^{\prime}}(E;\omega,{\bf q})$ contains the uncorrelated propagation of two particles, the product of the averaged retarded and advanced propagators, $G^{R}_{\mathbf{k}}(E_{+})$, and $G^{A}_{\mathbf{k}+\mathbf{q}}(E_{-})$, $E_{\pm}=E\pm\omega/2$, and a disorder-induced vertex correction
@@
where $L^{RA}$ is the electron-hole irreducible vertex, electron-hole self-energy. The electron-hole vertex, with all its variables, is important for the microscopic description. The solution of the problem of a particle moving in the infinite lattice with random atomic potentials is complete if we know the one-particle self-energy $\Sigma^{R}_{\mathbf{k}}(E)$ and the electron-hole irreducible vertex $L^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$.
The variables of the electron-hole vertex comprise microscopic, dynamical variables, such as wave vectors $\mathbf{k}$ and $\mathbf{k}^{\prime}$, as well as macroscopic, conserved quantities: Fermi energy $E$, transfer energy $\omega$, and transfer wave vector $\mathbf{q}$. The microscopic variables cannot be neglected in the microscopic theory of Anderson localization, but they affect global, measurable quantities, such as conductivity or diffusion, only in an averaged way. The fundamental function for the determination of the measurable transport properties is the electron-hole correlation function, defined as
@@
We use the Bethe-Salpeter equation (7) to reach the dominant contribution to the electron-hole correlation function in the low-energy limit $\omega,q\to 0$. We can determine the eigenvectors of the symmetrized version of the integral kernel in this equation. That is, we find normalized vectors $\phi^{S}_{\mathbf{k}}(E)$ obeying the following equations
@@
We assumed that the low-energy limit is dominated and controlled by the irreducible vertex $L^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ in the denominator $\mathcal{D}^{S}(E;\omega,\mathbf{q})$ without changing the eigenvector $\phi^{S}_{\mathbf{k}}(E)$. The eigenvector $\left|\phi\right\rangle$ with the maximal eigenvalue $S(E)=1$ then determines the low-energy asymptotics of the electron-hole correlation function. It is
@@
It is easy to see that the electron-hole correlation function contains a pole: $\mathcal{D}(E;0,\mathbf{0})=0$. It is essential, however, to know how the pole is reached when $\omega,q\to 0$.

### v4

@@ Abstract
The existence of Anderson localization, characterized by vanishing diffusion due to strong disorder, has been demonstrated in numerous ways. A systematic approach based on the Anderson quantum model of the Fermi gas in random lattices that can describe both diffusive and localized regimes has not yet been fully established. We build on a recent publication [1] and present a microscopic theory of disordered electrons that covers both the metallic phase with extended Bloch waves and the localized phase, where a propagating particle forms a quantum bound state with the hole left behind at the origin. The general theory provides a framework for constructing controlled approximations to one- and two-particle Green functions that satisfy the necessary conservation laws and causality requirements across the full range of disorder strength. It is used explicitly to derive a local, mean-field-like approximation for the two-particle irreducible vertices, enabling quantitative analysis of the solution’s dynamic properties in both metallic and localized phases, including critical behavior at the mobility edge. A new instability line for the dynamical electron-hole correlation function of the metallic phase is introduced.
@@ I Introduction
Anderson formulated a simple model of the propagation of a quantum particle in a random lattice [2]. His solution, in which diffusion is suppressed over long distances when disorder is sufficiently strong, is known as Anderson localization and sparked tremendous interest in the physics community. Early research focused on locating the mobility edge between extended and localized states of the Fermi gas in models of disordered alloys [3, 4, 5, 6, 7, 8]. The vanishing of diffusion was attributed to a zero of the static electric conductivity. Based on qualitative considerations and experimental evidence, Mott suggested a scenario of a discontinuous metal-insulator transition at the mobility edge with a minimal metallic conductivity [5, 9]. Symmetry arguments and the apparent short-range behavior of the averaged resolvent led to a conjecture that the critical behavior of interacting matrices, nonlinear $\sigma$-model, may simulate the mobility edge behavior in dimensions $2+\epsilon$ [10, 11, 12].
@@
The analogy between the Anderson localization transition and the critical behavior of phase transitions in statistical mechanics was supported by scaling arguments in Ref. [13]. Since then, the focus has shifted from the original Anderson quantum model of disordered non-interacting electrons to classical statistical mechanical models, allowing the use of established techniques for studying the critical behavior of interacting systems, such as scaling properties and universality classes [14]. The models differed in the universality class [11] or the underlying lattice [7]. In particular, a solution on the Bethe lattice with infinite coordination number was considered a mean-field solution of the Anderson localization transition. Unlike regular lattices, the critical behavior at the localization transition was found to be exponential [15, 16]. The exponential critical behavior was later shown to be an artifact of the Bethe lattice [17]. Moreover, the mean-field solution on the Bethe lattice was later shown to miss the correct high-dimensional limit of the Cayley tree [18, 19]. The Bethe lattice is the infinite limit of a regular random graph [20], and the concept of Anderson localization was then extended to random graphs [21, 22, 23, 24].
@@
Random matrix models in low spatial dimensions, $d\leq 2$, were known to exhibit a localized spectrum [25]. The existence of the Anderson localization transition and the mobility edge in dimensions $d>2$ was proved rigorously [26]. Numerical simulations in $d=3$ complemented the mathematical proof [27, 28]. The concept of Anderson localization has since broadened beyond its original formulation in charge diffusion and has been extended to light propagation [29, 30], the localization of thermal phonons [31], and the Anderson transition in cold atoms [32].
@@
All these approaches focused on the critical behavior of Anderson localization in the insulating phase. They use either localized states or a diverging scale, namely the localization length. They are inapplicable in the metallic phase, where no evident diverging scale exists. Instead, the limit of the static conductivity to zero has been used to indicate the Anderson localization transition. To find a theory of the Anderson localization transition approached from the metallic phase, the diagrammatic expansion in the Anderson model of disordered electrons, together with Green functions, was used to derive a self-consistent theory for the diffusion constant that turns zero at the localization transition [33, 34, 35, 36, 37]. It has been assumed to be an effective mean-field-like theory, for which, however, systematic extensions have not yet been found.
@@
Over the past six decades, we have gained significant insights into the Anderson localization transition and its associated critical behavior in the localized phase, characterized by a diverging localization length. However, not all questions about the origin of the localized phase have been answered. First, the critical scale is defined only in the localized phase. Second, no order parameter analogous to those in statistical mechanics has been identified. Third, the microscopic origin of the localized states, as described by the Anderson model, remains unexplained in existing approaches.
@@
The path to a microscopic theory of Anderson localization is to use the original model of a Fermi gas on a random lattice and solve it in a controlled manner. This is achieved by diagrammatically expanding the random potential in the averaged one- and two-particle resolvents. The well-established local mean-field theory for the electronic properties of random alloys, the coherent-potential approximation (CPA) [38, 39], does not lead to Anderson localization. A controlled way to go beyond the local approximation is to take the limit where the spatial dimension tends to infinity. The first fully self-consistent perturbation theory for single-electron properties, exact in infinite spatial dimensions, was shown to be equivalent to CPA [40, 41]. It misses vertex corrections to the Drude conductivity and is unfit to describe Anderson localization [42, 43]. Vertex corrections that diminish the electric conductivity are caused by backscattering and non-local correlations missed by the CPA [44, 45].
@@
The charge diffusion and its vanishing depend on how far a particle can propagate in a random environment. This information is not contained in the one-particle self-energy. It is the electron-hole correlation function, derived from the two-particle Green’s function, that carries information about the origin of electron propagation. It is determined from a Bethe-Salpeter equation with a two-particle irreducible vertex. The Anderson localization transition is signaled by a bifurcation point at which a new solution for the two-particle vertex emerges. Self-consistent nonlinear equations for the two-particle vertex must be used to see the bifurcation in the two-particle vertex [46]. We earlier used the parquet construction, combining Bethe-Salpeter equations for nonlocal irreducible vertices in the electron-hole and electron-electron scattering channels, to introduce two-particle self-consistency [47, 48]. This approach is similar to Refs. [33, 34, 37], which replaced the full two-particle self-consistency of the parquet equations for the irreducible vertices with a self-consistent equation for the dynamical diffusion coefficient. We identified a critical point that resembles the Anderson localization transition. Our solution, however, seemed unphysical since it did not obey the necessary Ward identity between the self-energy and the irreducible electron-hole vertex of Ref. [34]. This Ward identity guarantees the correct low-energy asymptotics of the diffusion pole in the density response function of the metallic phase. The dynamical vertex in the perturbation expansion is causal, so the related self-energy has the demanded analytic properties. We showed that microscopic causality is incompatible with the macroscopic Ward identity [49, 50]. The observed macroscopic quantities, the electron-hole correlation function and conductivity, must conform to the conservation laws. We showed how a conserving vertex complying with the desired Ward identity is obtained from the dynamical one of the perturbation theory in Ref. [51]. This addition enabled us to establish a controlled method for relating the outcome of microscopic perturbation theory to macroscopic observables. It opened a route towards a consistent microscopic theory of Anderson localization.
@@
I took the first step toward a fully controllable microscopic theory of Anderson localization in a recent publication [1]. A local approximation to the two-particle irreducible vertex in the parquet equations led to a critical point in the metallic phase of the Anderson model of disordered electrons. The critical point was identified with the Anderson localization transition. The diverging scale at the transition point was identified as the frequency derivative of the dynamical electron-hole irreducible vertex, or the dynamical conductivity. It was found to match the divergence of the localization length on the other side of the Anderson localization transition. The most fundamental conclusion of this paper was that the localized state is a quantum bound state of the propagating electron and the hole left at the beginning of its motion.
@@
This paper builds on Ref. [1] to present a fully controllable microscopic theory of Anderson localization based on self-consistent nonlinear equations for two-particle irreducible vertices, derived from the parquet theory that interconnects multiple scattering in the electron-hole, diffuson, and electron-electron, cooperon channels. We show that, to reach the critical point signaling a transition to the localized phase, two-particle functions must contain more information than that provided by the self-energy. The localized phase contains quantum bound states characterized by a new order parameter that induces a gap in the two-particle vertex function. Because there is no condensation of the bound states, no gap appears in the one-particle self-energy. We present details of the localized bound states, determine the threshold energy for bond breaking, and describe how the localized state converts into a metal. We also discuss the relationship between the microscopic dynamical quantities derived from perturbation theory and the macroscopic observables that obey conservation laws. We further specify the extent to which the Ward identities are valid in the metallic and localized phases. We demonstrate that linear-response theory breaks down in the localized phase and that static theories cannot fully describe Anderson localization. Finally, we discuss the relationship between the dynamical conductivity, the electron-hole correlation function, and the complex diffusion function in both metallic and localized phases.
The paper is structured as follows. The basic definitions, exact relations, and restrictions required to obtain a consistent solution of the Anderson model for disordered electrons in terms of the averaged one-particle and two-particle Green functions are summarized in Sec. II. The important pieces of the derivation of a consistent mean-field approximation that yields the Anderson localization transition, presented in earlier publications (Refs. [46, 45, 47, 48, 51]), are summarized in Sec. III. The detailed dynamical properties of this solution and how to obtain measurable quantities near the mobility edge and in the localized phase are discussed in Sec. IV. The impact and consequences of the presented mean-field solution on understanding Anderson localization are discussed in Sec. V.
@@ II.1 Hamiltonian and averaged resolvent
We assume the validity of the ergodic hypothesis, which states that spatial averaging is equivalent to configurational averaging. Consequently, all physical quantities in the thermodynamic limit will be configurationally averaged and translationally invariant. The averaged functions of interest, which determine all relevant physical quantities, are the one-particle and two-particle resolvents in the complex energy plane. The matrix representation of the one-particle resolvent for complex energy $z=E+i\eta$ in direct space reads
@@
where $\epsilon(\mathbf{k})$ is the lattice dispersion relation. The second equality is the Dyson equation, which introduces the self-energy, $\Sigma_{\mathbf{k}}(z)$, that contains all information about the impact of the random potential on electron propagation. The averaged one-particle Green function determines the electronic structure and spectral properties of the Fermi gas in a random lattice. It contains no information about the extent of the configurationally dependent electron wave functions at the given energy, $z=E+i0^{+}$.
@@ II.2 Two-particle vertex and electron-hole correlation function
The extent of the eigenstates of the random Hamiltonian affects the two-particle resolvent or the averaged two-particle Green function. The two-particle Green function for non-interacting particles in a random potential generally has two complex energies $z_{1}$ and $z_{2}$.
@@
The behavior of the two-particle Green function with energies $z_{1}\equiv E_{+}=E+\omega/2+i0^{+}$ and $z_{2}\equiv E_{-}=E-\omega/2-i0^{+}$ is decisive in determining whether the eigenstates of the random Hamiltonian are localized or extended. Here $E$ is the Fermi energy and $\omega$ is the transfer energy between particle and hole. These energies characterize the electron-hole resolvent $\mathcal{G}^{RA}$, the wave-vector representation of which is
@@
with the Fermi energy $E$. The transfer energy $\omega$ and wave vector $\mathbf{q}$ are conserved during the scattering processes when the two-particle Green function is decomposed into elementary scattering events. The dynamical variables affected by the random potential are wave vectors $\mathbf{k}$ and $\mathbf{k}^{\prime}$.
The averaged two-particle Green function $\mathcal{G}^{RA}_{{\bf k}{\bf k}^{\prime}}(E;\omega,{\bf q})$ contains the uncorrelated propagation of two particles, the product of the averaged retarded and advanced propagators, $G^{R}_{\mathbf{k}}(E_{+})$ and $G^{A}_{\mathbf{k}+\mathbf{q}}(E_{-})$, with $E_{\pm}=E\pm\omega/2$, and a disorder-induced vertex correction
@@
where $L^{RA}$ is the electron-hole irreducible vertex and electron-hole self-energy. The electron-hole vertex, with all its variables, is important for the microscopic description. The problem of a particle moving in an infinite lattice with random atomic potentials is solved completely if we know the one-particle self-energy $\Sigma^{R}_{\mathbf{k}}(E)$ and the electron-hole irreducible vertex $L^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$.
The variables of the electron-hole vertex include microscopic, dynamical variables, such as wave vectors $\mathbf{k}$ and $\mathbf{k}^{\prime}$, as well as macroscopic, conserved quantities: Fermi energy $E$, transfer energy $\omega$, and transfer wave vector $\mathbf{q}$. The microscopic variables cannot be neglected in the microscopic theory of Anderson localization, but they affect global, measurable quantities, such as conductivity or diffusion, only through averaging. The fundamental function for determining the measurable transport properties is the electron-hole correlation function, defined as
@@
We use the Bethe-Salpeter equation (7) to obtain the dominant contribution to the electron-hole correlation function in the low-energy limit $\omega,q\to 0$. We determine the eigenvectors of the symmetrized version of the integral kernel in this equation. That is, we find normalized vectors $\phi^{S}_{\mathbf{k}}(E)$ that satisfy the following equations
@@
We assumed that the low-energy limit is dominated and controlled by the irreducible vertex $L^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ in the denominator $\mathcal{D}^{S}(E;\omega,\mathbf{q})$, without changing the eigenvector $\phi^{S}_{\mathbf{k}}(E)$. The eigenvector $\left|\phi\right\rangle$ with the maximal eigenvalue $S(E)=1$ then determines the low-energy asymptotics of the electron-hole correlation function. It is
@@
It is easy to see that the electron-hole correlation function has a pole: $\mathcal{D}(E;0,\mathbf{0})=0$. It is essential, however, to know how the pole is approached as $\omega,q\to 0$.